OpenStudy (anonymous):

The minute hand of a clock is 6 inches long and moves from 12 to 4 o'clock. How far does the tip of the minute hand move? Express your answer in terms of pi and then round to two decimal places.

OpenStudy (anonymous):

u need to convert polar coordinates into cartesian system coordinates

OpenStudy (anonymous):

ur initial polar coordinates are (6,90deg.)

OpenStudy (anonymous):

final are (6 ,120 deg)

OpenStudy (anonymous):

i really dont understand that

OpenStudy (anonymous):

\[distance traveled by tip of minute hand \in one revolution or 360 degrees=2*\pi r\] \[distance traveled by tip of minute hand \in 120 degrees=\frac{ 2*pir }{360 }*120\] r=6" calculate.

OpenStudy (anonymous):

\[=\frac{ 2*\pi*6*120 }{360 }=4 \pi \]

OpenStudy (anonymous):

\[4 \pi =4*3.14=12.56" \]